# Graph Reinforcement Learning for Combinatorial Optimization: A Survey and Unifying Perspective

### v1

@@ 2.5.3 Policy Iteration Methods
Let us look at some concrete examples of algorithms. Dynamic Programming (DP), which is model-based, applies the PI scheme as described above. It is one of the earliest solutions developed for MDPs (Bellman, 1957). However, since it involves updating the value of every state in the entire state space, it is computationally intensive.
The Q-learning (Watkins & Dayan, 1992) algorithm is an off-policy TD method that follows the GPI blueprint. Unlike DP, it only requires samples of interactions with the environment. It is proven to converge to the optimal value functions and policy in the tabular case with discrete actions, so long as, in all the states, all actions have a non-zero probability of being sampled (Watkins & Dayan, 1992). The agent updates its estimates according to: {align} Q(s,a) ←Q(s,a) + α( r + γmax_a’ ∈A(s’)Q(s’,a’) - Q(s,a) )
@@ 2.5.4 Learning a Policy Directly
A possible downside of BC is that it violates the i.i.d. assumption and may suffer from compounding errors as soon as the policy makes a mistake and trajectories diverge from those seen during training. Follow-up works considered reducing the impact of this issue by enabling the agent to query the expert for more data (Ross et al., 2011) or using Generative Adversarial Networks (Ho & Ermon, 2016). The imitated policies can be used by themselves, or as an initialization for RL (Levine & Koltun, 2013; You et al., 2018a). IL has found success for robotics tasks such as autonomous vehicle navigation (Pomerleau, 1991; Codevilla et al., 2018) and flying drones (Silver et al., 2008).
This concludes our discussion of the ways in which a policy may be learned from data. Before moving on to describing model-based approaches for search and decision-time planning, let us take the opportunity to mention the relationship between the RL algorithms discussed up to this point and the graph representation learning methods covered in Section 2.6.2. Namely, the set of model weights $\Theta$ of a deep graph embedding approach serve as an effective basis for parameterizing a policy $\pi_{\Theta}$ or an approximation $\hat{Q}_{\Theta}$ of the true action-value function. For updating the parameters $\Theta$ using gradient descent, instead of using cross-entropy or Mean Squared Error losses with respect to the labels of “ground truth” examples as one would in a Supervised Learning setting, one computes losses with respect to learning targets that are defined by the RL algorithm.
@@ 2.5.5 Search and Decision-Time Planning Methods
Recall the fact that, in the case of model-based methods, we have access to the transition function $P$ and reward function $R$. This means that an agent does not necessarily need to interact with the world and learn directly through experience. Instead, the agent may use the model in order to plan the best course of action, and subsequently execute the carefully thought-out plan in the environment. Furthermore, decision-time planning concerns itself with constructing a plan starting from the current state that the agent finds itself in, rather than devise a policy for the entire state space.
@@
It is worth first discussing two of the simpler search techniques that often serve as reference points. Exhaustive Search (ES) refers to expanding all possible paths in the MDP from the current state and picking the best trajectory (i.e., the trajectory yielding the highest return). It is typically impractical due to the computational and memory requirements, but can serve as a useful benchmark on small problems. Greedy Search creates, at each step, a search tree of shallow depth rooted at the current state. It subsequently picks the best child node, and repeats the search with the child node at the root until a terminal state is reached. It is less resource-intensive than ES, but may also lead to short-sighted decisions. Despite this, it is commonly used in practice to good effect in a variety of problems (Cormen et al., 2022).
In many applications, however, the branching factor $b$ and depth $d$ of the search tree make it impossible to explore all paths, or even perform greedy search. There exist proven ways of reducing this space, such as alpha-beta pruning. However, its worst-case performance is still $\mathcal{O}(b^{d})$ (Russell & Norvig, 2020, Chapter 5.3). A different approach to breaking the curse of dimensionality is to use random (also called Monte Carlo) rollouts: to estimate the goodness of a position, run random simulations from a tree node until reaching a terminal state (Abramson, 1990; Tesauro & Galperin, 1997).
@@ 2.6 Artificial Neural Networks on Graphs
In this section, we discuss how one can represent graphs and tame their discrete structure for ML tasks. We begin with a broad overview of graph representation learning and traditional approaches. We then cover GNNs and related “deep” approaches for embedding graphs.
@@ 2.6.1 Graph Representation Learning
Suppose we wish to represent a graph as a vectorial input $\mathbf{x}$ that can be fed to a ML model, such as an MLP. An obvious choice is to take the adjacency matrix $\mathbf{A}$ and apply vectorization in order to transform it to a column vector. In doing so, however, two challenges become apparent. One is the fact that by relabeling the nodes in the graph according to a permutation, the input vector is modified substantially, and may result in a very different output when fed to a model, even though the structure has remained identical. Furthermore, if a new node joins the network, our previous model is no longer applicable by default since the shape of the input vector has changed. How might we address this, and design a better graph representation?
Such questions are studied in the field of graph representation learning (Hamilton et al., 2017a). Broadly speaking, the field is concerned with learning a mapping that translates the discrete structure of graphs into vectorial representations with which downstream ML approaches can work effectively. Work in this area is focused on deriving vectorial embeddings for a node, a subgraph, or an entire graph. A distinction can be drawn between shallow and deep embedding methods (Hamilton et al., 2017a). The former category consists of manually-designed approaches such as those using local node statistics, characteristic graph matrices, or graph kernels. In the latter category, methods typically use a deep neural network trained with gradient descent, and learn the representation in a data-driven fashion.
@@ 2.6.2 Deep Graph Embedding Methods
Broadly speaking, deep embedding methods rely on the idea of neighborhood aggregation: the representation of a node is determined in several rounds of aggregating the embeddings and features of its neighbors, to which a non-linear activation function is applied.
The learning architectures are usually parameter-sharing: the manner of performing the aggregation and the corresponding weights are the same for the entire network. Another important aspect for deep embedding methods is that they can incorporate task-specific supervision: the loss corresponding to the decoder can be swapped for, e.g., cross-entropy loss in the case of classification tasks. A final distinction in this family of approaches is the way subgraph embeddings are constructed: variants include performing a sum or mean over node embeddings in a subgraph (Duvenaud et al., 2015; Dai et al., 2016), introducing a dummy node (Li et al., 2017), using layers that perform clustering (Defferrard et al., 2016), or learning the hierarchical structure end-to-end (Ying et al., 2018).
Several works have proposed increasingly feasible versions of convolutional filters on graphs, based primarily on spectral properties (Bruna et al., 2014; Henaff et al., 2015; Defferrard et al., 2016; Kipf & Welling, 2017) or their approximations. An alternative line of work is based on message passing on graphs (Sperduti & Starita, 1997; Scarselli et al., 2009) as a means of deriving vectorial embeddings. Both Message Passing Neural Networks (Gilmer et al., 2017) and Graph Networks (Battaglia et al., 2018) are attempts to unify related methods in this space, abstracting the commonalities of existing approaches with a set of primitive functions. The term “Graph Neural Network” is used in the literature, rather loosely, as an umbrella term to mean a deep embedding method.
@@
structure2vec (S2V) (Dai et al., 2016) is one of the earlier GNN variants, and it is inspired by probabilistic graphical models (Koller & Friedman, 2009). The core idea is to interpret each node in the graph as a latent variable in a graphical model, and to run inference procedures similar to mean field inference (Wainwright & Jordan, 2008) and loopy belief propagation (Pearl, 1988) to derive vectorial embeddings. Additionally, the approach replaces the “traditional” probabilistic operations (sum, product, and renormalization) used in the inference procedures with nonlinear functions (namely, neural networks), yielding flexibility in the learned representation. It was shown to perform well for classification and regression in comparison to other graph kernels, as well as to be able to scale to medium-sized graphs representing chemical compounds and proteins.
@@ 3 Graph Structure Optimization
Given natural limitations in scope, we only consider works that (1) use a graph representation of the problem; and (2) train a policy for structure optimization via RL. Let us briefly discuss a strand of work that is related yet out-of-scope. Several works in the ML literature have considered the generation of graphs with similar properties to a provided dataset. This is typically performed using a deep generative model, and may be seen as an ML-based alternative for the classic graph generative models such as that of Barabási & Albert (1999). These works primarily use datasets of examples of the final graph (i.e, the “finished product’)’ and not of intermediate ones, which, in a sense, corresponds to the steps of the generation process itself. They additionally require large collections of related examples, which may not always be available depending on the domain.
In this area, works that use an auto-regressive model (such as LSTM or GRU) resemble MDP formulations; decisions such as the addition of an edge can be treated as a token in a sequence to be learned by the model. Some notable works in this area are the technique proposed by Li et al. (2018), GraphRNN (You et al., 2018b), and the Graph Recurrent Attention Network (Liao et al., 2019). Other types of generative models, such as Variational Autoencoders and Generative Adversarial Networks, were also applied for generating molecules (Kusner et al., 2017; Guimaraes et al., 2018; De Cao & Kipf, 2018; Jin et al., 2018).
@@
Figure 2: High-level illustration of how Graph Structure Optimization problems are approached with RL. The MDP starts from a (possibly empty) initial graph $G$ with an objective function value $\mathcal{F}(G)$. The topology of the graph is modified incrementally (e.g., via edge additions and removals) until a termination condition such as the exhaustion of a modification budget is met. The goal is for the objective function value $\mathcal{F}(G^{*})$ of the resulting graph $G^{*}$ to be maximally increased relative to the starting point.
Table 1: High-level summary of works that use Graph Reinforcement Learning for structure optimization.
Problem
Structural Modification
Objective Function
Base RL Algorithm Function Approximation Citation
Attacking GNNs
Edge addition and removal
Misclassification rate of GCN
DQN S2V (Dai et al., 2018)
Edge addition for newly injected nodes
Misclassification rate of GCN
DQN S2V (Sun et al., 2020)
Degree-preserving edge rewiring
Misclassification rate of GCN
DQN S2V (Ma et al., 2021)
Network Design
Edge addition
Resilience
DQN S2V (Darvariu et al., 2021a)
Edge addition with spatial restrictions
Resilience, Efficiency
UCT (Darvariu et al., 2023a)
Edge rewiring
Degree entropy, Maximum Entropy Random Walk
DQN S2V (Doorman et al., 2022)
Degree-preserving edge rewiring
Resilience, Efficiency
PPO FireGNN (GIN-based) (Yang et al., 2023b)
Edge addition
Recovering implicit objective
MaxEnt Inverse RL MPNN (Trivedi et al., 2020)
Causal Discovery
One-shot graph generation
Bayesian Information Criterion
Actor-Critic Encoder-Decoder (Zhu et al., 2020)
Node ordering
Bayesian Information Criterion
Actor-Critic Encoder-Decoder (Wang et al., 2021)
Node ordering
Bayesian Information Criterion
DQN Encoder-Decoder (Yang et al., 2023a)
Edge addition
Bayesian Information Criterion
UCT (Darvariu et al., 2023b)
Molecule Optimization
Edge addition
Drug-likeness, Synthetic accessibility
PPO GCN (You et al., 2018a)
Node addition, edge addition and removal
Drug-likeness, Similarity to prior molecule
DQN Molecular fingerprint (Zhou et al., 2019)
@@ 3.2 Network Design
There is a long tradition of using graphs to represent and analyze the properties of infrastructure networks such as power grids, computer networks, and metro transportation systems. Their construction (or design) for optimizing a given objective, on which there has been comparatively less work, has typically been approached using (meta)heuristic methods (Beygelzimer et al., 2005; Schneider et al., 2011; Wang et al., 2014).
Darvariu et al. (2021a) approached the problem of constructing graphs from the ground up or modifying an existing graph so as to optimize a global objective function. The considered objectives capture the resilience (also called robustness) of the network, quantified by the fraction of the nodes that needs to be removed before breakdown when the graph is exposed to random failures (Cohen et al., 2000) and targeted attacks (Cohen et al., 2001). The work is motivated by the importance of modern infrastructure networks such as power grids. The MDP formulation, which also breaks down the edge addition into two actions, is able to account for the exclusion of already-existing edges, leading to strictly better performance than RL-S2V. The authors showed that the proposed RNet-DQN method is able to learn better strategies for improving network robustness than prior heuristics, and can obtain generalization to larger networks.

### v2

@@ 2.3.3 Policy Iteration Methods
Let us look at some concrete examples of algorithms. Dynamic Programming (DP) applies the PI scheme as described above. It is one of the earliest solutions developed for MDPs (Bellman, 1957). It leverages the principle that sequential decision-making problems can be broken down into subproblems. The optimal solutions to the subproblems, once found, can be recursively combined to solve the original problem much more efficiently compared to algorithms that do not exploit this structure. Beyond an algorithmic paradigm, DP is also an exact method for combinatorial optimization as defined in Section 2.1, which highlights the shared foundation of RL and classic combinatorial optimization methods.
Since DP involves updating the value of every state, it is computationally intensive, and, for this reason, it is not appropriate for large problem instances. Additionally, it requires full knowledge of the transition and reward functions. Taken together, these characteristics limit its applicability, motivating the development of RL algorithms that do not require the exploration of the entirety of the state space and are able to learn directly from samples of interactions with the environment.
The Q-learning (Watkins & Dayan, 1992) algorithm is an off-policy TD method that follows the GPI blueprint. It is proven to converge to the optimal value functions and policy in the tabular case with discrete actions, so long as, in all the states, all actions have a non-zero probability of being sampled (Watkins & Dayan, 1992). The agent updates its estimates according to: {align} Q(s,a) ←Q(s,a) + α( r + γmax_a’ ∈A(s’)Q(s’,a’) - Q(s,a) )
@@ 2.3.5 Search and Decision-Time Planning Methods
Recall the fact that, in the case of model-based methods, we have access to the (possibly estimated) transition function $P$ and (possibly estimated) reward function $R$. This means that an agent does not necessarily need to interact with the world and learn directly through experience. Instead, the agent may use the model in order to plan the best course of action, and subsequently execute the carefully thought-out plan in the environment. Furthermore, decision-time planning concerns itself with constructing a plan starting from the current state that the agent finds itself in, rather than devise a policy for the entire state space.
@@
In many applications, however, the branching factor $b$ and depth $d$ of the search tree make it impossible to explore all paths, or even perform a Greedy Search of depth $1$. There exist proven ways of reducing this space, such as alpha-beta pruning. However, its worst-case performance is still $\mathcal{O}(b^{d})$ (Russell & Norvig, 2020, Chapter 5.3). A different approach to breaking the curse of dimensionality is to use random (also called Monte Carlo) rollouts: to estimate the goodness of a position, run random simulations from a tree node until reaching a terminal state (Abramson, 1990; Tesauro & Galperin, 1997).
@@ 2.4 Artificial Neural Networks on Graphs
In this section, we discuss how one can represent graphs and tame their discrete structure for combinatorial optimization tasks. We begin with a broad overview of graph representation learning and traditional approaches. We then cover GNNs and related “deep” approaches for embedding graphs. We refer the interested reader to (Hamilton, 2020) for an introduction to graph representation learning, which covers these techniques in more detail. Furthermore, the work of (Cappart et al., 2023) discusses the use of such methods, particularly GNNs, for combinatorial optimization.
@@ 2.4.1 Graph Representation Learning
Suppose we wish to represent a graph as a vector input $\mathbf{x}$ that can be fed to a ML model, such as an MLP. An obvious choice is to take the adjacency matrix $\mathbf{A}$ and apply vectorization in order to transform it to a column vector. In doing so, however, two challenges become apparent. One is the fact that by relabeling the nodes in the graph according to a permutation, the input vector is modified substantially, and may result in a very different output when fed to a model, even though the structure has remained identical. Furthermore, if a new node joins the network, our previous model is no longer applicable by default since the shape of the input vector has changed. How might we address this, and design a better graph representation?
Such questions are studied in the field of graph representation learning (Hamilton et al., 2017a). Broadly speaking, the field is concerned with learning a mapping that translates the discrete structure of graphs into vector representations with which downstream ML approaches can work effectively. Work in this area is focused on deriving vector embeddings for a node, a subgraph, or an entire graph. A distinction can be drawn between shallow and deep embedding methods (Hamilton et al., 2017a). The former category consists of manually-designed approaches such as those using local node statistics, characteristic graph matrices, or graph kernels. In the latter category, methods typically use a deep neural network trained with gradient descent, and learn the representation in a data-driven fashion.
@@ 2.4.2 Deep Graph Embedding Methods
Figure 2: Illustration of the neighborhood aggregation principle. To determine the features of a node, those of its neighbours are aggregated using learnable parametrizations, to which an activation function is applied. While the particulars depend on the architecture, many deep embedding methods follow this blueprint.
Broadly speaking, deep embedding methods rely on the idea of neighborhood aggregation: the representation of a node is determined in several rounds of aggregating the embeddings and features of its neighbors, to which a non-linear activation function is applied. This principle is illustrated in Figure 2.
The learning architectures are usually parameter-sharing: the manner of performing the aggregation and the corresponding weights are the same for the entire network. Another important aspect for deep embedding methods is that they can incorporate task-specific supervision: the loss corresponding to the decoder can be swapped for, e.g., cross-entropy loss in the case of classification tasks.
Many deep representation learning techniques can be used to construct representations for a graph or subgraph starting from the embeddings of the nodes. Several variants that have been considered in the literature include: performing a sum or mean over node embeddings in a subgraph (Duvenaud et al., 2015; Dai et al., 2016), introducing a dummy node (Li et al., 2017), using layers that perform clustering (Defferrard et al., 2016), or learning the hierarchical structure end-to-end (Ying et al., 2018).
Several works have proposed increasingly feasible versions of convolutional filters on graphs, based primarily on spectral properties (Bruna et al., 2014; Henaff et al., 2015; Defferrard et al., 2016; Kipf & Welling, 2017) or their approximations. An alternative line of work is based on message passing on graphs (Sperduti & Starita, 1997; Scarselli et al., 2009) as a means of deriving vector embeddings. Both Message Passing Neural Networks (Gilmer et al., 2017) and Graph Networks (Battaglia et al., 2018) are attempts to unify related methods in this space, abstracting the commonalities of existing approaches with a set of primitive functions. The term “Graph Neural Network” is used in the literature, rather loosely, as an umbrella term to mean a deep embedding method.
@@
structure2vec (S2V) (Dai et al., 2016) is one of the earlier GNN variants, and it is inspired by probabilistic graphical models (Koller & Friedman, 2009). The core idea is to interpret each node in the graph as a latent variable in a graphical model, and to run inference procedures similar to mean field inference (Wainwright & Jordan, 2008) and loopy belief propagation (Pearl, 1988) to derive vector embeddings. Additionally, the approach replaces the “traditional” probabilistic operations (sum, product, and renormalization) used in the inference procedures with nonlinear functions (namely, neural networks), yielding flexibility in the learned representation. It was shown to perform well for classification and regression in comparison to other graph kernels, as well as to be able to scale to medium-sized graphs representing chemical compounds and proteins.
@@
2.5 Connections Between RL and Graph Representation Learning
Given the RL algorithms and graph representation learning techniques covered up to this point, let us take the opportunity to discuss the relationships between them in the context of the Graph RL framework. Learning techniques designed for operating on graphs are commonly used as function approximators as part of the RL algorithms. However, not all works covered by the survey employ function approximation – notably, some of the works that use MCTS-based approaches do not. Furthermore, out of the works employing function approximation, a few opt for different learning architectures through problem-specific justifications or experimental validation. Generic architectures such as MLPs, LSTMs, and Transformers are sometimes employed. Nevertheless, as discussed in our inclusion criteria in Section 1, we require that works formulate graph combinatorial optimization problems as MDPs, but we do not restrict ourselves to particular forms of function approximation.
Therefore, the distinction between Graph RL and graph representation learning is indeed blurred, as the latter is often (but not always) a component of the former. When Graph RL solutions do leverage graph representation learning for function approxmation, they are a means of obtaining generalization to unseen instances, as well as for scaling up to larger instances in terms of number of nodes and edges and the complexity of the considered process taking place on the graph. Furthermore, we note the following difference between the use of graph representation learning techniques as function approximators in this context versus the typical benchmarks considered in the graph learning literature. Graph RL methods are constructive and goal-driven, which allows for flexibility in discovering embeddings relevant for the objective function to be optimized without a fine-grained supervision signal. Instead, standard graph learning benchmarks rely on supervised learning and the availability of granular examples (real or synthetic data).
Recall that we denote the set of all weights of a model as $\Theta$. Function approximation methods based on graph representation learning techniques have been used to parameterize the policy $\pi_{\Theta}$ (Ma et al., 2021; Yang et al., 2023b) as well as the action-value function $\hat{Q}_{\Theta}$ (Dai et al., 2018; Zhou et al., 2019). The surveyed works that use function approximation mostly employ either of the two in isolation, or combine them in Actor-Critic style architectures (Meirom et al., 2021). As previously discussed, the reward function $R$ can generally be computed analytically or estimated via sampling for the considered problems. An exception in this sense is the work on molecular optimization of (You et al., 2018a), who used a Graph Convolutional Network as a discriminator trained on example molecules to provide part of the reward signal. None of the works we surveyed learn the transition model $P$. However, as we have argued, a fully accurate transition model is known for many of the problems.
It is worth mentioning another recent line of work that connects RL and GNNs. While none of the surveyed techniques employ this connection, they are nevertheless relevant in a broader sense. Specifically, this literature draws a connection between GNNs and Dynamic Programming. In this work, a graph representation is constructed in which nodes are possible states in the MDP, and edges correspond to transitions determined by the actions. With this framing, GNNs can be seen as “executing neurally”, in an approximate fashion, the steps of DP algorithms. Thus, this can be applied to standard RL problems.
In this line of work, Xu et al. (2020) empirically showed that computations carried out by GNNs are aligned with those performed by the Bellman-Ford algorithm, which uses DP. Dudzik & Veličković (2022) later established this connection formally for a broader class of DP algorithms. Deac et al. (2020) empirically demonstrated that, with sufficiently granular supervision, GNNs can accurately model the Value Iteration algorithm for DP. Lastly, Deac et al. (2021) took this idea further by integrating learning a model and planning in latent space. We note that the graph structure used in these works is a higher-level representation of the MDP itself and not of a combinatorial optimization problem defined on the graph – and indeed, such techniques have been applied to continuous control tasks. As the scalability of such techniques is currently limited, they have not yet been applied to combinatorial optimization problems.
@@ 3 Graph Structure Optimization
Given the scope of the survey as defined in Section 1, we exclude several works which, at a high level, share notable similarities. Let us briefly discuss two strands of work that are related yet out-of-scope. Firstly, several works in the ML literature have considered the generation of graphs with similar properties to a provided dataset. This is typically performed using a deep generative model, and may be seen as an ML-based alternative for the classic graph generative models such as that of Barabási & Albert (1999). These works primarily use datasets of examples of the final graph (i.e, the “finished product’)’ and not of intermediate ones, which, in a sense, corresponds to the steps of the generation process itself. They additionally require large collections of related examples, which may not always be available depending on the domain. In this area, works that use an auto-regressive model (such as LSTM or GRU) resemble MDP formulations; decisions such as the addition of an edge can be treated as a token in a sequence to be learned by the model. Some notable works in this area are the technique proposed by Li et al. (2018), GraphRNN (You et al., 2018b), and the Graph Recurrent Attention Network (Liao et al., 2019). Other types of generative models, such as Variational Autoencoders and Generative Adversarial Networks, were also applied for generating molecules (Kusner et al., 2017; Guimaraes et al., 2018; De Cao & Kipf, 2018; Jin et al., 2018). Secondly, optimizing the structure of graphs has also been considered by works that apply RL agents to physical construction environments, such as stacking blocks (Hamrick et al., 2018; Bapst et al., 2019). These works use a graph representation in which objects correspond to blocks, and edges indicate physical relationships between them. A Graph Network (Battaglia et al., 2018) is used to predict $Q$-values for each edge. Results in this area indicate that agents equipped with the inductive bias derived from this representation can attain better performance for physical construction problems. However, as these works do not treat a discrete optimization problem but instead use the graph representation as part of a more complex algorithm for a different task, we do not cover them in detail.
@@
Figure 3: High-level illustration of how Graph Structure Optimization problems are approached with RL. The MDP starts from a (possibly empty) initial graph $G$ with an objective function value $\mathcal{F}(G)$. The topology of the graph is modified incrementally (e.g., via edge additions and removals) until a termination condition, such as the exhaustion of a modification budget is met. The goal is for the objective function value $\mathcal{F}(G^{*})$ of the resulting graph $G^{*}$ to be maximally increased relative to the starting point.
@@ 3.1 Attacking Graph Neural Networks
Traditional approaches for adversarial attacks can be divided into white-box methods (which assume access to the internals of the classifier, such as gradients) and black-box methods (which do not require this privileged information). For the latter setting, metaheuristics that perform gradient-free optimization, such as evolutionary algorithms, can be leveraged (Su et al., 2019). Additionally, an attack based on random edge rewiring can also be considered; this can induce substantial classification errors despite its apparent simplicity (Dai et al., 2018). Recent works have therefore considered using RL to improve on the performance of these attack methods in the absence of access to the internal state of classifiers.
@@ 3.2 Network Design
There is a long tradition of using graphs to represent and analyze the properties of infrastructure networks such as power grids, computer networks, and metro transportation systems. Their construction (or design) for optimizing a given objective, on which there has been comparatively less work, has typically been approached using (meta)heuristic methods. Several notable examples of traditional methods include the work of Beygelzimer et al. (2005), who considered heuristics for edge addition and rewiring based on random selection and node degrees. Schneider et al. (2011) proposed a method based on the simulated annealing metaheuristic for rewiring infrastructure networks. Lastly, heuristics that leverage the spectrum of the graph Laplacian (Wang & Van Mieghem, 2008; Wang et al., 2014) to guide modifications have also been considered.
In the space of RL methods, Darvariu et al. (2021a) approached the problem of constructing graphs from the ground up or modifying an existing graph so as to optimize a global objective function. The considered objectives capture the resilience (also called robustness) of the network, quantified by the fraction of the nodes that needs to be removed before breakdown when random failures (Cohen et al., 2000) and targeted attacks (Cohen et al., 2001) occur. The work is motivated by the importance of modern infrastructure networks such as power grids. The MDP formulation, which also breaks down the edge addition into two actions, is able to account for the exclusion of already-existing edges, leading to strictly better performance than RL-S2V. The authors showed that the proposed RNet-DQN method is able to learn better strategies for improving network robustness than prior heuristics, and can obtain generalization to larger networks.